# Brenda Prallon22footnotemark: 2    Chen Qiu22footnotemark: 2    Yiwei Sun22footnotemark: 2

### v1

@@ 5.1 Mobile Financial Services in Bangladesh
Figure 5(b) reports the time needed to implement the synthetic purposive sampling approach of Egami and Lee (2024). This was done by using the spsR package with the option to use the Gurobi solver on the background. We consider it to be fast, taking less than $40$ seconds for $k=10$. However, the synthetic purposive sampling problem can be formulated as a quadratic mixed integer program, while $k$-median is linear. This is evident from the graph; the $k$-median problem is solved almost instantly for every $k$ up to $10$.
@@ 5.2 Multi-Country Survey Experiments in Europe
Our second application revisits a multi-country survey experiment originally conducted and analyzed in Naumann et al. (2018) and discussed in Egami and Lee (2024). The question of interest is whether native-born inhabitants of a particular country are more supportive of immigration depending on whether the potential migrants are high-skilled or low-skilled. Naumann et al. (2018) carried out a survey experiment in $15$ European countries: Austria, Belgium, Czech Republic, Denmark, Finland, France, Germany, Ireland, Netherlands, Norway, Slovenia, Spain, Sweden, Switzerland, and the United Kingdom. Respondents were native-born individuals and were randomly assigned to report their attitudes towards either high-skilled (“treatment”) or low-skilled (“control”) immigrants.
While the experiments have already been conducted and outcomes of each experiment are available for all countries, we consider the situation of a researcher that can only conduct $k=6$ experiments; here the choice of $k=6$ mirrors Egami and Lee (2024). We let all $15$ countries be both potential experimental and policy-relevant sites. That is, $\mathcal{S}_{E}=\mathcal{S}_{P}$, and card($\mathcal{S}_{E}$) = card($\mathcal{S}_{P}$) = 15.  [^16] A potential situation we have in mind is one of a researcher who would like to give policy advice to the policy makers in these countries on whether to initiate an immigration reform that could favor either high-skilled or low-skilled immigrants. The researcher knows that policy makers are interested in voters’ attitudes towards these types of reforms. We assume that the researcher is only able to experiment in a subset of countries due to administrative or budget constraints, and that he/she needs to extrapolate voters’ attitudes of the other policy-relevant sites based on the experimental estimates. The researcher needs to decide whether to recommend the implementation of an immigration reform directed to either high-skilled or low-skilled immigrants.
@@
Figure 7(a) also shows that three selected countries in the $k$-median problem (Switzerland, Czech Republic, and Spain) are not connected to any of the other policy-relevant sites. This suggests that these countries are selected because no other country provides a close enough match for them. In contrast, in Figure 7(b), these countries (Switzerland, Czech Republic, and Spain) receive positive weights from at least five other countries.
@@
Approximation Error in Theorem 1: Because experimental estimates and corresponding standard errors are available for all $15$ countries, we can compute the approximation error in Theorem 1. Define the relative approximation error of the $k$-median approximation to minimax regret as
$\frac{\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)}{\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)}.$ (23)
@@
$\max\left\{0,1-\frac{B\cdot\sigma_{E}\cdot\frac{\min\{\#\left(\mathcal{S}_{E}\cap\mathcal{S}_{P}\right),k\}}{\#\mathcal{S}_{P}}}{\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)}\right\},$ (24)
@@
$1+\frac{B\cdot\sigma_{E}\cdot\frac{\min\{\#\left(\mathcal{S}_{E}\cap\mathcal{S}_{P}\right),k\}}{\#\mathcal{S}_{P}}}{\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|\right)},$ (25)
@@
Figure 8 displays these bounds for values of $k\in\{1,\ldots,10\}$. As $k$ increases, the approximation becomes worse; as mentioned before, this is driven by ignoring the experimental sites themselves in bounding regret. In this application, the $k$-median solution still works relatively well when $k=6$, with a relative error of about ±$\pm 13\%$. As $k$ increases to $10$, the relative error is about ±$\pm 50\%$. [^18] For these computations, we picked $C$ to be the smallest Lipschitz constant needed to capture the heterogeneity of estimated treatment effects in the data. That is, we pick $C$ as $\max_{s,s^{\prime}\in\mathcal{S}_{E}\cup\mathcal{S}_{P}}\frac{|\hat{\tau}_{s}-\hat{\tau}_{s^{\prime}}|}{\|X_{s}-X_{s^{\prime}}\|},$ (26) where $\hat{\tau}_{s}$ denote the estimated treatment effects for site $s$. It turns out that this $C$ is comfortably “large” in the sense of Theorem 1. We provide additional details in Appendix B.3.
@@ 6.1 Fixed Costs of Experimentation
$\mathcal{W}_{c}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\left(\sum_{s\in\mathcal{S}_{P}}\tau(X_{s})\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]-\sum_{s\in\mathscr{S}}c_{s}\right).$ (27)
@@
$\mathcal{R}_{c}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}}c_{s}+\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\tau(X_{s})\left(\mathbf{1}\{\tau(X_{s})\geq 0\}-\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right).$ (28)
Under the assumptions of Theorem 1, it is possible to show (by extending Lemma 1 and 2 to account for the fixed costs of experimentation) that
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}^{1/2}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}_{c}(T,\mathscr{S},\tau)\right)$
@@
The problem in (29) is the metric uncapacitated $k$-facility location problem that was discussed in Remark 7. This is a common extension of the $k$-median problem where there is a fixed cost of opening each facility. The problem can also be formulated as a linear integer program, namely
$\min_{\{y_{i},x_{i,j}\}_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}}$ $\left(\sum_{i\in\mathcal{S}_{E}}y_{i}c^{*}_{i}+\sum_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}x_{i,j}\cdot c(j,i)\right)$ (30)
such that $\sum_{i\in\mathcal{S}_{E}}x_{i,j}=1,\;\;\;\;\;\forall j\in\mathcal{S}_{P},$ (31)
$\sum_{i\in\mathcal{S}_{E}}y_{i}\leq k,$ (32)
$0\leq x_{i,j}\leq y_{i},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P},$ (33)
$y_{i}\in\{0,1\},x_{i,j}\in\{0,1\},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}.$ (34)
Just as before, the choice variable $y_{i}$ indicates whether facility $i$ is open, and $x_{i,j}$ indicates whether client $j$ is assigned to facility $i$. Constraints are imposed to ensure that each client $j$ is assigned to at least one facility, that there are no more than $k$ facilities in total, and that a client can only be assigned to an open facility. The connection cost between a facility $i$ and a client $j$ is still given by $\|X_{i}-X_{j}\|$, but now there is a fixed cost of opening a facility $i$ given by $c^{*}_{i}\equiv 2c_{i}/C$. This means that—in contrast to the case in which there is no fixed cost—choosing the optimal sites to maximize external validity now requires knowledge of the Lipschitz constant $C$ (as this constant appears explicitly on the fixed cost $c^{*}_{i}$). This is simply because, as we have shown, the scale of regret absent costs of experimentation depends on $C$. Consequently, every time a new site is considered for experimentation, there is a trade-off between its contribution to reduce regret and the fixed cost of experimentation. We note that when $C$ is large, the fixed costs in the objective function become negligible, and the solution of the uncapacitated facility location problem can be approximated by the solution of the $k$-median problem.
@@ 6.2 Random Selection of Experimental Sites
So far, our analysis focused on minimizing the worst-case (welfare-based) regret among all purposive sampling schemes that select at most $k$ sites. That is, we excluded randomized (including nonuniformly randomized) sampling schemes. This is an important limitation—a widespread view among experimenters is that “the external validity of randomized evaluations for a given population (say, the population of a country) would be maximized by randomly selecting sites and, within these sites, by randomly selecting treatment and comparison groups” (Duflo et al., 2007, p. 3953).
@@
In our setting, whether the decision maker will want to randomize crucially depends on timing assumptions; that is, the moment in the game in which nature can move to harm the decision maker. Consider first the case in which an adversarial nature may harm the policy maker by choosing $\tau$ only after seeing the realization of $\mathscr{S^{*}}$ (and knowing $\mathcal{T}_{\mathscr{S}^{*}}$). Then, the risk of using treatment rule $T\in\mathcal{T}_{\mathscr{S}^{*}}$ is
$\mathcal{R}(T,\mathscr{S}^{*},\tau),$
@@
$\inf_{p\in\Delta(\mathcal{A}(k))}\left(\sum_{\mathscr{S}^{*}\in\mathcal{A}(k)}p(\mathscr{S}^{*})\cdot V(\mathscr{S}^{*})\right).$
@@
That said, the timing that seems more in line with Wald’s (1950) original application of the minimax principle is likely one in which the policy maker commits to both a randomized sampling scheme and a set of contingent (on sampling scheme) decision rules and nature adversarially picks $\tau$ before any randomization was realized. We next briefly discuss this possibility.
[…]
To see that solving this problem can quickly become very hard, consider now the same example, except that the experimental sites coincide with the policy sites at $\{1,2\}$. Then we can find the MMR optimal combination of purposive sampling scheme and treatment assignment rule, and we can also verify that randomized site selection will strictly reduce worst-case regret. However, we are unable to characterize the exact solution for this, still extremely structured, case. We leave further exploration of randomized site selection, including the potentially fruitful investigation of “good enough” (in terms of regret) such schemes, for future research.
@@ 6.3 Other Restrictions on Treatment Heterogeneity
With Assumption 1, the effect of the policy of interest is a Lipschitz function of observed covariates with respect to the Euclidean distance. This means that the effect of the policy of interest in two sites with the same covariates is assumed to be the same (in other words, there is no unobserved treatment effect heterogeneity). In this section, we show that our main result can also accommodate other general measures of the distance between covariates (thus allowing for other functional classes), as well as some forms of unobserved treatment effect heterogeneity (provided $\mathcal{S}_{E}\cap\mathcal{S}_{P}=\emptyset$).
@@ 6.3.1 Other Distance Measure Based on Observed Covariates
To see how our results can allow for other types of distances, let $m(x,x^{\prime}):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ be a measure of distance (metric) between $x,x^{\prime}\in\mathbb{R}^{d}$. [^20] That is, we assume that $m(x,x^{\prime})>0$ for any $x\neq x^{\prime}$, and that for any $x$ we have $m(x,x)=0$. We also assume that the function is symmetric, in that $m(x,x^{\prime})=m(x^{\prime},x)$. And finally, we assume that $m$ satisfies the triangle inequality $m(x,x^{\prime})\leq m(x,x^{\prime\prime})+m(x^{\prime},x^{\prime\prime})$ for any $x,x^{\prime},x^{\prime\prime}$. Also see Dudley (2002, p.20). Then, Assumption 1 may be modified as:
@@
For example, $m(x,x^{\prime})=\left(\left(x-x^{\prime}\right)^{\top}\left(x-x^{\prime}\right)\right)^{1/2}$ is the Euclidean distance in Assumption 1 and $m(x,x^{\prime})=\left(\left(x-x^{\prime}\right)^{\top}W\left(x-x^{\prime}\right)\right)^{1/2}$ for some positive definite matrix $W$ is a $W$-weighted Euclidean distance. Choosing $m(x,x^{\prime})=\left[\tilde{m}(x,x^{\prime})\right]^{\alpha}$ for some $\alpha\in(0,1)$ and some distance measure $\tilde{m}(\cdot,\cdot)$ also effectively allows us to model $\tau$ as a Hölder continuous function of order $\alpha$ (Dudley, 2002, p. 56). With Assumptions 2 and 3, $N_{\mathscr{S}}(s)$ is understood to be a nearest neighbor in $\mathscr{S}$ measured in terms of metric $m(\cdot,\cdot)$ that appears in Assumption 3. Moreover, $\text{Lip}_{C}(\mathbb{R}^{d})$ now stands for the space of all Lipschitz functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ with constant $C$, but in terms of metric $m(\cdot,\cdot)$. We note that we are slightly abusing notation, because we are not using $m$ to index the nearest-neighbor function $N_{\mathscr{S}}(s)$ or the Lipschitz functional class $\text{Lip}_{C}(\mathbb{R}^{d})$, although their respective definitions explicitly on the choice of metric $m(\cdot,\cdot)$. Under a general metric $m(\cdot,\cdot)$, the $k$-median problem is simply modified as:
@@
[…]

### v2

@@ 5.1 Mobile Financial Services in Bangladesh
Figure 5(b) reports the time needed to implement the synthetic purposive sampling approach of Egami and Lee (2024). This was done by using the spsR package with the option to use the Gurobi solver on the background. We consider it to be fast, taking less than $40$ seconds for $k=10$. However, while synthetic purposive sampling can be formulated as a quadratic mixed integer program, $k$-median is linear and correspondingly faster to solve; indeed, solutions were almost instant for every $k$ up to $10$.
@@ 5.2 Multi-Country Survey Experiments in Europe
Our second application revisits a multi-country survey experiment originally conducted and analyzed in Naumann et al. (2018) and discussed in Egami and Lee (2024). The question of interest is whether native-born inhabitants of a particular country are more supportive of immigration depending on whether the potential migrants are high-skilled or low-skilled. Naumann et al. (2018) carried out a survey experiment in $15$ European countries listed in Figure 7. Respondents were native-born individuals and were randomly assigned to report their attitudes towards either high-skilled (“treatment”) or low-skilled (“control”) immigrants.
While the experiments have already been conducted and outcomes of each experiment are available for all countries, we follow Egami and Lee (2024) and consider the situation of a researcher that can only conduct $k=6$ experiments. We let all $15$ countries be both potential experimental and policy-relevant sites. That is, $\mathcal{S}_{E}=\mathcal{S}_{P}$, and $\mathcal{S}_{E}=\#\mathcal{S}_{P}=15$. [^17] A potential situation we have in mind is one of a researcher who would like to give policy advice to the policymakers in these countries on whether to initiate an immigration reform that could favor either high-skilled or low-skilled immigrants. The researcher knows that policymakers are interested in voters’ attitudes towards these types of reforms. We assume that the researcher is only able to experiment in a subset of countries due to administrative or budget constraints, and that he/she needs to extrapolate voters’ attitudes of the other policy-relevant sites based on the experimental estimates. The researcher needs to decide whether to recommend the implementation of an immigration reform directed to either high-skilled or low-skilled immigrants.
@@
Figure 7(a) also shows that three selected countries in the $k$-median problem (Switzerland, Czech Republic, and Spain) are not connected to any of the other policy-relevant sites, suggesting that they were selected because no other country provides a close enough match for themselves. In contrast, in Figure 7(b), these countries receive positive weights from at least five other countries.
@@
Approximation Error in Theorem 1: Because experimental estimates and corresponding standard errors are available for all $15$ countries, we can compute the approximation error in Theorem 1. Define the relative approximation error of the $k$-median approximation to MMR as
$\frac{\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}(T,\mathscr{S},\tau)\right)}{\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\inf_{\mathscr{S}\in\mathcal{A}(k)}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|}.$ (12)
@@
$\max\left\{0,1-\frac{B\cdot\sigma_{E}\cdot\frac{\min\{\#\left(\mathcal{S}_{E}\cap\mathcal{S}_{P}\right),k\}}{\#\mathcal{S}_{P}}}{\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\inf_{\mathscr{S}\in\mathcal{A}(k)}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|}\right\},$
@@
$1+\frac{B\cdot\sigma_{E}\cdot\frac{\min\{\#\left(\mathcal{S}_{E}\cap\mathcal{S}_{P}\right),k\}}{\#\mathcal{S}_{P}}}{\frac{C}{2}\frac{1}{\#\mathcal{S}_{P}}\inf_{\mathscr{S}\in\mathcal{A}(k)}\sum_{s\in\mathcal{S}_{P}\backslash\mathscr{S}}\|X_{s}-X_{N_{\mathscr{S}}(s)}\|},$
@@
Figure 8 displays these bounds for values of $k\in\{1,\ldots,10\}$. As $k$ increases, the approximation becomes worse; as mentioned before, this is driven by ignoring the experimental sites themselves in bounding regret. In this application, the $k$-median solution still works relatively well when $k=6$, with a relative error of about $\pm 13\%$. As $k$ increases to $10$, the relative error is about $\pm 50\%$. [^19] For these computations, we picked $C$ to be the smallest Lipschitz constant needed to capture the heterogeneity of estimated treatment effects in the data. That is, we pick $C$ as $\max_{s,s^{\prime}\in\mathcal{S}_{E}\cup\mathcal{S}_{P}}\frac{|\widehat{\tau}_{s}-\widehat{\tau}_{s^{\prime}}|}{\|X_{s}-X_{s^{\prime}}\|},$ (13) where $\widehat{\tau}_{s}$ denote the estimated treatment effects for site $s$. It turns out that this $C$ is comfortably “large” in the sense of Theorem 1. We provide additional details in Appendix B.3.
@@ 6.1 Fixed Costs of Experimentation
$\mathcal{W}_{c}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\left(\sum_{s\in\mathcal{S}_{P}}\tau(X_{s})\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]-\sum_{s\in\mathscr{S}}c_{s}\right).$
@@
$\mathcal{R}_{c}(T,\mathscr{S},\tau):=\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathscr{S}}c_{s}+\frac{1}{\#\mathcal{S}_{P}}\sum_{s\in\mathcal{S}_{P}}\tau(X_{s})\left(\mathbf{1}\{\tau(X_{s})\geq 0\}-\mathbb{E}_{\tau_{\mathscr{S}}}[T_{s}(\widehat{\tau}_{\mathscr{S}})]\right).$
Under the assumptions of Theorem 1, one can show that
$\inf_{\mathscr{S}\in\mathcal{A}(k)}\left(\inf_{T\in\mathcal{T}_{\mathscr{S}}}\sup_{\tau\in\textrm{Lip}_{C}(\mathbb{R}^{d})}\mathcal{R}_{c}(T,\mathscr{S},\tau)\right)$
@@
The problem in (14) is the metric uncapacitated $k$-facility location problem discussed in Remark 2. This is a common extension of the $k$-median problem where there is a fixed cost of opening each facility. It can also be formulated as a linear integer program, namely
$\min_{\{y_{i},x_{i,j}\}_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}}$ $\left(\sum_{i\in\mathcal{S}_{E}}y_{i}c^{*}_{i}+\sum_{i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}}x_{i,j}\cdot c(j,i)\right)$
such that $\sum_{i\in\mathcal{S}_{E}}x_{i,j}=1,\;\;\;\;\;\forall j\in\mathcal{S}_{P},$
$\sum_{i\in\mathcal{S}_{E}}y_{i}\leq k,$
$0\leq x_{i,j}\leq y_{i},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P},$
$y_{i}\in\{0,1\},x_{i,j}\in\{0,1\},\;\;\;\;\;i\in\mathcal{S}_{E},j\in\mathcal{S}_{P}.$
Just as before, the choice variable $y_{i}$ indicates whether facility $i$ is open, and $x_{i,j}$ indicates whether client $j$ is assigned to facility $i$. Constraints are imposed to ensure that each client $j$ is assigned to at least one facility, that there are no more than $k$ facilities in total, and that a client can only be assigned to an open facility. The connection cost between a facility $i$ and a client $j$ is still given by $\|X_{i}-X_{j}\|$, but now there is a fixed cost $c^{*}_{i}\equiv 2c_{i}/C$ of opening a facility $i$. Choosing the optimal sites now requires knowledge of the Lipschitz constant $C$ as this constant appears explicitly in the fixed cost $c^{*}_{i}$. Consequently, every time a new site is considered for experimentation, there is a trade-off between its contribution to reduce regret and the fixed cost of experimentation. When $C$ is large, the fixed costs in the objective function become negligible, and the solution of the uncapacitated facility location problem can be approximated by the solution of the $k$-median problem.
@@ 6.2 Random Selection of Experimental Sites
So far, our analysis focused on minimizing the worst-case (welfare-based) regret among all purposive sampling schemes that select at most $k$ sites. That is, we excluded randomized (including nonuniformly randomized) sampling schemes. This is an important limitation—a widespread view among experimenters is that “the external validity of randomized evaluations for a given population (say, the population of a country) would be maximized by randomly selecting sites and, within these sites, by randomly selecting treatment and comparison groups” (Duflo et al., 2007, p. 3953). A similar point is made in (List, 2024, p. 493): “if the researcher chooses locations at random in an initial stage of the experimental design, the this will lead to generalizable results across all potential locations.”
@@
In our setting, whether the decision maker will want to randomize crucially depends on timing assumptions; that is, the moment in the statistical game in which nature can move. Consider first the case in which an adversarial nature may choose $\tau$ after seeing the realization of $\mathscr{S^{*}}$ (and knowing $\mathcal{T}_{\mathscr{S}^{*}}$). Then, the risk of using treatment rule $T\in\mathcal{T}_{\mathscr{S}^{*}}$ is
$\mathcal{R}(T,\mathscr{S}^{*},\tau)$
@@
$\inf_{p\in\Delta(\mathcal{A}(k))}\sum_{\mathscr{S}^{*}\in\mathcal{A}(k)}p(\mathscr{S}^{*})\cdot V(\mathscr{S}^{*}).$
@@
That said, the timing that seems more in line with Wald’s (1950) original application of the minimax principle is one in which the policymaker commits to both a randomized sampling scheme and a set of contingent (on sampling scheme) decision rules and nature adversarially picks $\tau$ before any randomization was realized. We next briefly discuss this possibility.
To see that randomization might strictly speaking be optimal, consider a stylized example where $k=1$ and the covariates of each site are equal to its index: $S_{E}=\{1,4\}$, and $S_{P}=\{2,3\}$. For simplicity, suppose furthermore that $\widehat{\tau}_{s}=\tau_{s}$, i.e. there is no sampling uncertainty in the treatment effects. This example can be solved for those combinations of sampling scheme and treatment assignment rule that achieve exact MMR. As we formally show in Appendix B.2, the exact MMR attainable by purposive sampling equals $3C/4$, whereas the exact MMR with randomized sampling equals C/2C/2.202020The assumption of perfect signals is for simplicity. The example is rigged to resemble cases analyzed in Montiel Olea et al. (2025) and Stoye (2012), and we would accordingly be able to generalize it, but for our present purpose, solving for arbitrary sampling variances would only add tedium. Thus, in principle there can be a gain to randomized sampling.
To see that solving this problem can quickly become very hard, consider now the same example, except that the experimental sites coincide with the policy sites at $\{1,2\}$. Then we can find the MMR optimal combination of purposive sampling scheme and treatment assignment rule, and we can also verify that randomized site selection will strictly reduce worst-case regret. However, we are unable to characterize the exact solution for this, still extremely structured, case.
In related work, Fernández et al. (2024) further explores the use of the Hedge Algorithm for finding the minimax regret optimal randomized site selection. Their results suggest that even if randomization is allowed, it is possible that choosing the site that is most representative for the policy-relevant site could still be approximately minimax regret optimal. Moreover, in the application they consider, the approximately optimal randomized selection schemes are far from uniform sampling.
@@ 6.3.1 Other Distance Measure Based on Observed Covariates
To see how our results can allow for other types of distances, let $m(x,x^{\prime}):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ be a distance metric on $\mathbb{R}^{d}$. [^21] That is, we assume that $m(x,x^{\prime})>0$ for any $x\neq x^{\prime}$, and that for any $x$ we have $m(x,x)=0$. We also assume that the function is symmetric, in that $m(x,x^{\prime})=m(x^{\prime},x)$. And finally, we assume that $m$ satisfies the triangle inequality $m(x,x^{\prime})\leq m(x,x^{\prime\prime})+m(x^{\prime},x^{\prime\prime})$ for any $x,x^{\prime},x^{\prime\prime}$. Also see Dudley (2002, p.20). Then, Assumption 1 may be modified as:
@@
[…]
@@
Since the proof of Theorem 1 does not rely on any specific property of Euclidean distance (other than the fact that it is a distance), Theorem 1 still holds after replacing the Euclidean distance $\left\|\cdot$-$\cdot\right\|$ with an arbitrary metric $m(\cdot,\cdot)$.
A natural question that arises in light of the objective function in (15) is whether there is a weaker version of Assumption 3 where $m(\cdot,\cdot)$ is not necessarily a metric. Such a result could be helpful, for example, to provide a decision-theoretic foundation for use of the k-means algorithm for site selection, which would arise by setting $m(x,x^{\prime})=\|x-x^{\prime}\|^{2}$. Unfortunately, it is known that the squared Euclidean distance is not a metric (as it violates triangle inequality). Moreover, it is also known that if one insists in having a version of Assumption 3 with $m(x,x^{\prime})=\|x-x^{\prime}\|^{2}$, the only functions that satisfy this property in $\mathbb{R}^{d}$ are constant functions, which preclude treatment effect heterogeneity. Thus, while arbitrary Lipschitz-type conditions for general functions $m(\cdot,\cdot)$ are likely not useful, in Section 6.3.4 we discuss the extent to which we can generate results for a further generalization of Assumption 3 that considers a convex and centrosymmetric space of functions.
In practice, as with matching estimators, distance measures must weight covariates, but there is no obvious best way to do so, and the solution to the $k$-median problem will be sensitive to this choice. Following the literature (Imbens, 2004; Imbens and Wooldridge, 2009; Imbens and Rubin, 2015, Section 18.5), we think two metrics are reasonable: Mahalanobis metric, i.e., $m(x,x^{\prime})=\bigl{(}(x-x^{\prime})^{\top}\Sigma_{X}^{-1}(x-x^{\prime})\bigr{)}^{1/2}$, where $\Sigma_{X}$ is the covariance matrix of site-level covariates; or the modification thereof that sets all off-diagonal elements of $\Sigma_{X}^{-1}$ to zero. Relative to Euclidean distance, the Mahalanobis metric has the advantage of being invariant to changes in the scale of covariates (and more generally, to affine transformations of covariates). However, this metric is not invariant to other transformations and it is not usually recommended for categorical variables. For alternative options, in Section 6.3.3 below we also discuss how a structural model for treatment effects can be used naturally to define a metric $m(\cdot,\cdot)$ over covariates.